A manufacturer of fluorescent tubes claims that average life rime is more than 1750 hour, A sample of 400 fluorescent light bulbs produced mean life tone 1600 hours with a standard deviation of ISO hours. Does the claim of the company justified at 5% level of significance?


1
Expert's answer
2021-07-19T05:48:55-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq 1570"

"H_1:\mu>1570"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05, df=n-1"

"=400-1=399" degrees of freedom, and the critical value for a right-tailed test is "t_c=1.648682."

The rejection region for this right-tailed test is "R=\{t:T>1.648682\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{1600-1570}{150/\sqrt{400}}=4"

Since it is observed that "t=4>1.648682=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for right-tailed "\alpha=0.05, df=399, t=4" is

"p=0.000038," and since "p=0.000038<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"  is greater than "1570," at the "\alpha=0.05" significance level.